If 4.97 mol of C5H12 reacts with excess O2, how many moles of CO2 will be produced by the following combustion reaction?

C5H12+8O2⟶6H2O+5CO2
1
Expert's answer
2020-09-21T06:37:25-0400

From the equation it is seen that from one mol of C5H12 five moles of CO2 is formed. So, the number of moles of CO2 may be found from a proportion:

x=4.97×5/1=24.85(mol).